OpenStudy (anonymous):

Susan spends four hours a day,Sunday through Friday,surfing the internet.During the entire week,she spends an average of 5 hours per day surfing.On Saturday,how many hours does she spend on the internet?

OpenStudy (anonymous):

Could someone help me

OpenStudy (anonymous):

[(6)(4) + x] / 7 = 5 where x is the # of hours on Saturday. So, multiply both sides by 7 and then subtract 24 from both sides to get "x".

OpenStudy (anonymous):

Its a bit hard for me to understand

OpenStudy (anonymous):

I'll rewrite it using the equation tool so it's easier to see.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

\[\frac{ (6 \times 4) + x }{ 7 } = 5\]The "6" is from six days (not including Saturday). "x" is Saturday hours. When you add them all together for (6 x 4) + x you get the total # of hours of internet surfing for the week and now you divide by 7 to get the average per day, which we know to be 5 because we were given that information.

OpenStudy (anonymous):

When you multiply both sides by 7 you get: (6 x 4) + x = 5 x 7 = 35 "x" can be solved from this as it becomes: 24 + x = 35 Subtract 24 from each side: x = 11

OpenStudy (anonymous):

So the answer is 11

OpenStudy (anonymous):

Yes. x = 11. Nice working with you again and good luck in all of your studies! And thx for the recognition! @momof4

OpenStudy (anonymous):

Your welcome & thanks again